OpenStudy (anonymous):

If f(x) = −2x − 1 and g(x) = 6x − 1, find f[g(−2)]

OpenStudy (blurbendy):

g(-2) = 6(-2) - 1 g(-2) = -12 -1 g(-2) = -13 put -13 in for f(x)

OpenStudy (anonymous):

Is it 25?